OpenStudy (anonymous):

The measures of 6 of the interior angles of a heptagon are: 120°, 150°, 135°, 170°, 90°, and 125°. What is the measure of the largest exterior angle 90° 80° 75° 85°

OpenStudy (anonymous):

 substitute all the interior angles in place of n and get the largest exterior angle

OpenStudy (anonymous):

i dont think its that way

OpenStudy (anonymous):

isnt your answer 90

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

total mesure of haptagon is (n-1)180=6*180=1080